# Controlling dynamical systems into unseen target states using machine learning

### v1

@@ Abstract
We present a novel, model-free, and data-driven methodology for controlling complex dynamical systems into previously unseen target states, including those with significantly different and complex dynamics. Leveraging a parameter-aware realization of next-generation reservoir computing, our approach accurately predicts system behavior in unobserved parameter regimes, enabling control over transitions to arbitrary target states. Crucially, this includes states with dynamics that differ fundamentally from known regimes, such as shifts from periodic to intermittent or chaotic behavior. The method’s parameter-awareness facilitates non-stationary control, ensuring smooth transitions between states. By extending the applicability of machine learning-based control mechanisms to previously inaccessible target dynamics, this methodology opens the door to transformative new applications while maintaining exceptional efficiency. Our results highlight reservoir computing as a powerful alternative to traditional methods for dynamic system control.
@@ Introduction
Controlling dynamical systems is a fundamental challenge in engineering and science, crucial for maintaining stability, optimizing performance, and enabling adaptability across a broad spectrum of applications. From aerospace engineering, where precise control of aircraft and spacecraft is vital [1, 2], to industrial automation, where maintaining efficiency and safety in manufacturing processes is essential [3], the ability to effectively control dynamical systems is driving significant advancements among different fields [4, 5, 6, 7, 8]. In biomedical engineering, controlling physiological systems and biomedical devices improved patient care[9, 10, 11]. In the realm of energy systems, it enabled the efficient operation of power grids and the integration and optimization of renewable energy sources [12, 13]. However, the increasing integration of distributed renewable energy sources poses challenges to the robustness of power grids. Consequently, it is imperative to develop new control strategies tailored for future micro grids to ensure stability and reliability [14]. Traditionally, control methods such as Proportional-Integral-Derivative (PID) controllers, state-space control, and model predictive control (MPC) have been employed to manage these systems. While effective in many cases, these approaches often rely on accurate mathematical models and can struggle with the complexities inherent in nonlinear, high-dimensional, and time-varying systems [15, 16, 17]. In recent years, the advent of machine learning introduced powerful new tools for predicting and controlling dynamical systems. Techniques such as deep learning, neural ordinary differential equations and reinforcement learning showed promise in modeling complex systems and developing control strategies [18, 19, 20]. However, these methods often require extensive computational resources and can be difficult to interpret and implement. Reservoir computing (RC) emerges as a promising candidate in this landscape, offering a simpler yet effective approach for handling the temporal dependencies and nonlinearities of dynamical systems. Reservoir computing leverages a fixed, high-dimensional dynamical system, the reservoir, to process temporal data. The key advantage of reservoir computing lies in its separation of dynamic processing, managed by the reservoir, from the training process, which optimizes only the readout layer. This significantly reduces the computing costs and makes it orders of magnitude more efficient than comparable algorithms[21, 22]. Recent advancements in reservoir computing led to the development of various reservoir architectures, enhancing their predictive capabilities and computational efficiency. Traditional reservoir computing employs a randomly generated network at its core [23], providing robust performance in learning and predicting temporal patterns. Minimal reservoir computing simplifies this approach significantly by using a block-diagonal matrix structure, which lowers computational complexity and minimizes training data requirements [24]. Additionally, the next-generation reservoir (NGRC) computing architecture, originating from nonlinear vector autoregression, utilizes a deterministic library of unique monomials derived from the input data and time shifts of it [25]. Studies comparing these architectures demonstrated their effectiveness across various applications. For instance, the prediction performance of these architectures on chaotic systems showed that minimal reservoir computing and next-generation reservoir computing outperforms traditional methods in scenarios with limited data availability, providing more accurate and reliable predictions [24]. Further research showed that next-generation reservoir computing can achieve comparable predictive performance to the already data-efficient traditional reservoir computing architectures while requiring only a fraction of the training data, about $10$ to $100$ times less [25, 26, 27]. These innovations make reservoir computing as a framework more flexible and effective, especially in scenarios with limited data and strong nonlinearity, opening up its use in various fields and problems. One field is the control of dynamical systems. In [28], traditional reservoir computing is used to control dynamical systems into arbitrary periodic, intermittent or chaotic target states without requiring knowledge of the underlying governing equations. To achieve successful control, the reservoir computer needs to accurately learn the statistical behavior of the target dynamic. Then, a control force based on the RC prediction and the current system state can be applied on the system, forcing it into the desired target state of the RC. In [27] the control performance of this approach was compared to that with NGRC at its core. While both methods demonstrate similar performance in reproducing the climate of the target states during control, NGRC outperformed RC by requiring approximately ten times less training data to achieve these results. Here we leverage this method to control dynamical systems into previously unseen target states by extrapolating these states from few training data samples using parameter-aware NGRC[29]. Further, the parameter-aware approach allows for non-stationary prediction. With that we introduce a novel non-stationary controlling method that ensure smooth transitioning from the current system state to the desired target state.
@@ Results
Prediction of unseen parameter states
@@
We assess the performance of the proposed method on the Lorenz system by varying the bifurcation parameter $\rho$ within the range previously analyzed for the prediction of unseen states using traditional reservoir computing [30]. For $\rho>100$, the system exhibits periodic behavior. As $\rho$ decreases below 100, it undergoes period-doubling bifurcations and transitions to chaotic dynamics. The bifurcation behavior is visualized in red within Fig. 1 $\mathbf{A}$. We use the parameter-aware NGRC approach, proposed in [29], to learn the Lorenz system using four periodic and stationary training data samples with parameters $\rho_{\text{train}}=100.0,100.1,100.2$ and $100.3$. After successful training we are able to extrapolate and interpolate the system behavior in unseen parameter regions of $\rho$, showing in Fig. 1 $\mathbf{B}$. For each parameter, we use a total data length of $5.000$ time steps for training the NGRC architecture. Compared to the $250.000$ training data points and additional $50.000$ synchronization data points per parameter used in [30], it becomes significant that the NGRC approach is highly data efficient when learning such dynamics.
Control into unseen parameter states
Accurate prediction of the dynamics of previously unseen states now allows for the application of the control mechanism described in [28, 27] to control the system dynamics toward these states. A control force is applied, defined as the product of a force scaling parameter $K$ and the difference between the system’s current state at $\rho_{\text{test}}$ and the predicted state of the target dynamics at $\rho_{\text{pred}}$. This control force directs the system’s dynamics at each time step to converge with the predicted dynamics. For example, the dynamic at $\rho_{\text{test}}=100.7$ can be controlled to align with any $\rho$ value within the analyzed bifurcation parameter range, as shown in Fig. 1 $\mathbf{C}$. This control mechanism is not limited to NGRC and is applicable to any method capable of accurately extrapolating and interpolating system dynamics.
Figure 1: Controlling dynamics into unseen states. A) Segments of the Lorenz bifurcation diagram (red). Four training samples and their bifurcation parameters are used for training (light green). B) Predicted bifurcation diagram for unseen parameter regions using the trained model. C) Controlled bifurcation diagram demonstrates the ability of the control mechanism to control system states into arbitrary and unseen target states using the models predictions.
Figure 2: Non-stationary transition duration influences system behavior. The orange line in (a) and (b) shows the parameter transition over time, while the red line represents the corresponding system dynamics from simulations. In (a), an abrupt parameter transition causes turbulent behavior after the transition, as also indicated by the increased running correlation dimension in the lower plot. In (b), a slow parameter transition leads to a smooth progression towards the target dynamics.
Non-stationary control
@@
The parameter-aware NGRC approach has been shown to predict non-stationary time series effectively [29]. This enables not only direct control of the trajectory towards the target system but also introduces non-stationary transitions by controlling the system from $\rho_{\text{test}}$ through intermediate $\rho$ values. Figure 2(a) demonstrates that abrupt changes in $\rho$ (e.g. $100.5$ to $99.5$) can result in turbulent [31] behavior post-transition, while Figure 2(b) shows smooth transitions to target dynamics with gradual parameter changes. An examination of the occurrence of turbulent behavior during different transition times is conducted, considering the potential adverse effects of turbulence in practical and technological applications. We compare the non-stationary simulation behavior with the non-stationary predictions of the parameter-aware NGRC model and develop a control mechanism that completely eliminates the occurrence of turbulent behavior, regardless of the transition time. We analyze how the system simulations as well as the NGRC predictions behave when they transition from a periodic orbit at $\rho_{\text{start}}=100.5$ to a chaotic regime at $\rho_{\text{target}}=99.5$. Here, the transitions are performed with linearly changing $\rho$ values. Depending on the transition time and the location of the trajectory when the parameter transition starts, turbulent behavior might occur. We utilize a running correlation dimension[32, 33] (rcd) approach as an operational measure to detect changes in the system behavior. As shown in Figure 2(a), a global maximum of the $rcd$ forms during the transition phase, while in Figure 2(b), the global maximum of the $rcd$ is defined by the target dynamics at $\rho_{\text{target}}$. Since the $rcd$ of the target dynamics at $\rho_{\text{target}}$ does not exceed 1.21, we define $rcd_{\text{limit}}=1.21$ as the $rcd$-cut-off value. If the $rcd$ values of the non-stationary simulations or predictions remains below this threshold over time, we assume that the trajectory is smoothly transitioning into the target dynamics. Conversely, for simulations or predictions where the $rcd$ values exceed this threshold, we assume that they exhibit turbulent behavior, at least to some extent. Starting from 100 random initial conditions at $\rho_{\text{start}}$, we determine for different transition times how likely it is that the trajectories exhibit turbulent behavior during the transition phase to $\rho_{\text{target}}$. In Figure 3, we show that these probabilities are strongly dependent on the transition time and are approaching zero for long enough transition periods. The probabilities observed in the simulations are reproduced by the NGRC predictions, underscoring that the NGRC approach learns both stationary and non-stationary behavior patterns from only a few stationary data samples. Based on these findings a control mechanism is derived that avoids any possible turbulent behavior during the parameter transitions. As NGRC allows for efficient prediction of multiple realizations of test trajectories, we exploit that for each transition time at least some predictions are showing smooth transition behavior without turbulence occurring. These predictions can then be used to control the system dynamic’s at $\rho_{\text{start}}$, firstly for some time at $\rho_{\text{start}}$ and than for the smooth transition to $\rho_{\text{target}}$. By applying this approach to all 100 initialized systems at $\rho_{\text{start}}$ for each transition time, the control mechanism reduces the turbulence probability to zero. In Fig. 4 one of the controlled trajectories for a abrupt transition with no transition time is shown. To evaluate the quality of the controlled dynamics we compare the statistical climate of the actual system at $\rho_{\text{target}}$ to that of the $100$ controlled systems for each transition period. Therefore we measure the spatial complexity as well as the temporal complexity for each controlled system by calculating its correlation dimension[33] and its largest Lyapunov exponent[34], respectively. For the actual system and each controlled system, the measurements were carried out over $25.000$ time steps and are displayed in Fig. 5. The correlation dimension of the actual system is $cd=1.1378$, reproduced with an average correlation dimension of $cd=1.1382$ across all controlled systems. Similarly, the largest Lyapunov exponent of the actual system is $ly=0.0735$, while the average largest Lyapunov exponent across all controlled systems is $ly=0.0628$. Both metrics, particularly the correlation dimension, indicate that the statistical climate of the actual system is effectively predicted following each transition period, despite the model being trained on periodic trajectories and is extrapolating the emerging chaotic system behavior. With that the method enables smooth control into unseen target states avoiding adverse turbulent behavior during transitions.
Figure 3: Control with smooth transitions into unseen target states. Starting from $100$ initial conditions, the probability of turbulent behavior for the simulation (red) and the model prediction (green) is shown in red and green respectively, given different transition periods and a parameter transitions from a period orbit at $\rho_{start}=100.5$ to chaotic one at $\rho_{target}=99.5$. If predictions that show no turbulent behavior are taken as control dynamics, the turbulence probability can be reduced to zero.
Figure 4: Smooth control into target state . In (a) the initial periodic system dynamics at $\rho_{start}=100.5$ without control is plotted. In (b), the trajectory is smoothly controlled into the chaotic target dynamics at $\rho_{target}=99.5$ using the proposed control mechanism. A smooth transition is provided even if an abrupt transition to the target parameter is specified.
@@
Figure 5: Statistical climate of controlled target states. Evaluation of the statistical climate of the $100$ controlled states at $\rho_{start}=99.5$ given each transition period shown in Fig. 3. In (a), the spatial complexity of the dynamics are measured using their correlation dimension. In (b), the temporal complexities are measured using the largest Lyapunov exponent.

### v2

@@ Abstract
We present a novel, model-free, and data-driven methodology for controlling complex dynamical systems into previously unseen target states, including those with significantly different and complex dynamics. Leveraging a parameter-aware realization of next-generation reservoir computing (NGRC), our approach accurately predicts system behavior in unobserved parameter regimes, enabling control over transitions to arbitrary target states utilizing a new prediction evaluation and selection scheme. Crucially, this includes states with dynamics that differ fundamentally from known regimes, such as shifts from periodic to intermittent or chaotic behavior. The method’s parameter awareness facilitates non-stationary control with which control scenarios are generated and evaluated on the basis of predefined control objective. In addition to proving the method for transient-free control to extrapolated chaotic target states over transition times, we demonstrate the method’s effectiveness on a nonlinear power system model. Our method successfully navigates transitions even in scenarios where system collapse is observed frequently, while ensuring fast transitions and avoiding prolonged transient behavior. By extending the applicability of machine learning-based control mechanisms to previously inaccessible target dynamics, the methodology opens the door to new control applications while maintaining exceptional efficiency.
@@ Introduction
Controlling dynamical systems is a fundamental challenge in engineering and science, crucial for maintaining stability, optimizing performance, and enabling adaptability across a broad spectrum of applications. From aerospace engineering, where precise control of aircraft and spacecraft is vital [46, 8], to industrial automation, where maintaining efficiency and safety in manufacturing processes is essential [31], the ability to effectively control dynamical systems is driving significant advancements among different fields [40, 37, 26, 11, 32]. In biomedical engineering, controlling physiological systems and biomedical devices improved patient care[10, 29, 20]. In the realm of energy systems, it enabled the efficient operation of power grids and the integration and optimization of renewable energy sources [6, 2]. However, the increasing integration of distributed renewable energy sources poses challenges to the robustness of power grids. Consequently, it is imperative to develop new control strategies tailored for future micro grids to ensure stability and reliability [45]. Traditionally, control methods such as Proportional-Integral-Derivative (PID) controllers, state-space control, and model predictive control (MPC) have been employed to manage these systems. While effective in many cases, these approaches often rely on accurate mathematical models and can struggle with the complexities inherent in nonlinear, high-dimensional, and time-varying systems [21, 1, 7]. In recent years, the advent of machine learning introduced powerful new tools for predicting and controlling dynamical systems. Techniques such as deep learning, neural ordinary differential equations and reinforcement learning showed promise in modeling complex systems and developing control strategies [36, 5, 9]. However, these methods often require extensive computational resources and can be difficult to interpret and implement. Reservoir computing (RC) emerges as a promising candidate in this landscape, offering a simpler yet effective approach for handling the temporal dependencies and nonlinearities of dynamical systems. Reservoir computing leverages a fixed, high-dimensional dynamical system, the reservoir, to process temporal data. The key advantage of reservoir computing lies in its separation of dynamic processing, managed by the reservoir, from the training process, which optimizes only the readout layer. This significantly reduces the computing costs and makes it orders of magnitude more efficient than comparable algorithms[43, 19]. Recent advancements in reservoir computing led to the development of various reservoir architectures, enhancing their predictive capabilities and computational efficiency. Traditional reservoir computing employs a randomly generated network at its core [22], providing robust performance in learning and predicting temporal patterns. Minimal reservoir computing simplifies this approach significantly by using a block-diagonal matrix structure, which lowers computational complexity and minimizes training data requirements [34]. Additionally, the next-generation reservoir (NGRC) computing architecture, originating from nonlinear vector autoregression, utilizes a deterministic library of unique monomials derived from the input data and time shifts of it [14]. Studies comparing these architectures demonstrated their effectiveness across various applications. For instance, the prediction performance of these architectures on chaotic systems showed that minimal reservoir computing and next-generation reservoir computing outperforms traditional methods in scenarios with limited data availability, providing more accurate and reliable predictions [34]. Further research showed that next-generation reservoir computing can achieve comparable predictive performance to the already data-efficient traditional reservoir computing architectures while requiring only a fraction of the training data, i.e. about $10$ to $100$ times less [14, 3, 17]. These innovations make reservoir computing as a framework more flexible and effective, especially in scenarios with limited data and strong nonlinearity, opening up its use in various fields and problems. One field is the control of dynamical systems [18, 17, 23, 24]. In [18], traditional reservoir computing is used to control dynamical systems into arbitrary periodic, intermittent or chaotic target states without requiring knowledge of the underlying governing equations. To achieve effective control, the reservoir computer must accurately capture the statistical behavior of the target dynamics. Subsequently, a control force—derived from the RC’s prediction and the current system state—is applied to control the system toward the desired target state. In [17], the performance of this approach was compared with one using NGRC at its core. Although both methods perform similarly in reproducing the statistical climate of the target states during control, NGRC outperformed the RC approach by requiring roughly ten times less training data. In this work, we now make use of the extrapolative capabilities of parameter-aware NGRC by learning a dynamical digital twin of the dynamical systems analyzed from few parametrized data sets[27]. We introduce a general prediction evaluation and selection scheme to generate control strategies that fulfill a predefined control objective, with which the dynamical systems are controlled in extrapolated and previously unseen target states. Thereby we demonstrate the use of the introduced control methodology in critical control scenarios, where previous control approaches would lead to transient behavior or even system collapse.
@@ Results
We evaluated the proposed control method on two dynamical systems with distinct objectives and application contexts. First, we control the Lorenz system from an initially periodic orbit to an extrapolated weakly chaotic state, imposing the requirement that no chaotic transients arise during the parameter transition, a phenomenon that is frequently observed in simulations of this parameter change but is consistently suppressed by our approach across a wide range of transition times. Second, we apply the controller to a power system model whose dynamics is governed by reactive power demand and are prone to rich chaos and even voltage collapse. Starting in a periodic regime, we execute an instantaneous parameter change to an extrapolated periodic window located just before the collapse threshold, achieving fast transitions to the target state while reliably preventing both collapse and the prolonged chaotic transients typically observed in simulations. In both studies, we train a parameter-aware NGRC on few parameterized data sets, allowing it to serve as a dynamical digital twin that efficiently generates non-stationary candidate trajectories toward extrapolated targets, which are evaluated and selected if they satisfy a desired control objective to become the control strategy. For the Lorenz system and its weakly chaotic target state, we utilize a running correlation dimension metric for the evaluation scheme. For the power system model and its initial periodic state and periodic target state, we evaluate the control scenarios generated by measuring the spectral entropy over time with the objective of minimizing the transition duration.
Lorenz system
@@
We test the method on the Lorenz system by varying the parameter $\rho$ in the vicinity of $\rho=100$. For $\rho>100$, the system exhibits periodic behavior. As $\rho$ decreases below 100, it undergoes period-doubling bifurcations and transitions to weakly chaotic dynamics. The bifurcation behavior is colored red in Fig. LABEL:fig:turbulence_probability_prediction $\mathbf{A}$ and has previously been analyzed for the prediction of unseen states using traditional reservoir computing [25]. We use parameter-aware NGRC [27] to learn the Lorenz system using four periodic and stationary training data samples with parameters $\rho_{\text{train}}=100.0,100.1,100.2$ and $100.3$. After successful training, the method can extrapolate and interpolate the behavior of the system in the unseen parameter regions of $\rho$, as shown in Fig. LABEL:fig:turbulence_probability_prediction $\mathbf{B}$. For each parameter, a total data length of $5.000$ time steps is sufficient to train the NGRC. We use the trained NGRC as the basis for the control approach and leverage its dynamical similarity to the actual dynamical system to generate scenarios for the control strategy. First, we test for the simple case of stationary target states as a control objective, inter- and extrapolated by our parameter-aware approach. A control force is applied, defined as the product of a control parameter $K$ and the difference between the current state of the system at $\rho_{\text{test}}$ and the predicted state of the target dynamics at $\rho_{\text{pred}}$. The control force directs the system’s dynamics at each time step to converge with the predicted dynamics. For example, the dynamic at $\rho_{\text{test}}=100.7$ can be controlled to align with any $\rho$ value within the range of the analyzed bifurcation diagram, as shown in Fig. LABEL:fig:turbulence_probability_prediction $\mathbf{C}$.
@@
Subsequently, we test the control approach for non-stationary dynamics. As shown in Fig. LABEL:fig:turbulence $\mathbf{A}$, abrupt changes in $\rho$ (e.g. from $100.5$ to $99.5$) can lead to transient behavior in the system dynamics post-transition [35], whereas the gradual parameter change in Fig. LABEL:fig:turbulence $\mathbf{B}$ results in a smooth transition. This behavior is further analyzed in the right plot of Fig. 1, where we analyze how the system simulations behave when they transition from a periodic orbit at $\rho_{\text{start}}=100.5$ to a weakly chaotic regime at $\rho_{\text{target}}=99.5$. Here, the transitions are performed with linearly changing $\rho$ values. Depending on the transition time and the location of the trajectory when the parameter transition starts, transient behavior might occur. We capture the transient behavior with a running correlation dimension (rcd) [30, 16] approach (detailed in Methods), by defining that if the rcd becomes larger during the transition than that of the weakly chaotic target state, transient behavior occurred. For each analyzed transition time, we apply this measure on 100 randomly initialized Lorenz systems to define a probability of how likely it is for the transition time that the system exhibiting transient dynamics. Now, we leverage this measure to generate control strategies for each transition time to avoid transients in the system during the parameter change. Therefore we employ the parameter-aware NGRC to predict from different initial conditions non-stationary control scenarios, inducing transitional periods from $\rho_{\text{start}}$ through intermediate $\rho$ values to the extrapolated target state at $\rho_{\text{target}}$, and select those scenarios, where the rcd stays below the extrapolated maximum of the weakly chaotic target state during the transition as the control strategy. With this approach we are able to reduce the transient probability for all analyzed transition periods quantitatively to zero, shown in Fig. 1 $\mathbf{D}$. In Fig. 1 $\mathbf{A}$ the initial system dynamic of the Lorenz system at $\rho_{\text{start}}$ is shown, together with a controlled trajectories for an abrupt transition (Fig. 1 $\mathbf{B}$) and its rcd (Fig. 1 $\mathbf{C}$). The weakly chaotic state is visible in changes in amplitudes of the oscillations after the parameter change. We evaluate the quality of the controlled dynamics by comparing the statistical climate of the actual system at $\rho_{\text{target}}$ to that of the $100$ controlled systems for each transition period. Therefore we measure the spatial complexity as well as the temporal complexity for each controlled system by calculating its correlation dimension[16] and its largest Lyapunov exponent[41], respectively. For the actual system and each controlled system, the measurements were carried out over $25.000$ time steps and are displayed in Fig. 2. The correlation dimension of the actual system is $cd=1.1378$, reproduced with an average correlation dimension of $cd=1.1382$ across all controlled systems. Similarly, the largest Lyapunov exponent of the actual system is $ly=0.0735$, while the average largest Lyapunov exponent across all controlled systems is $ly=0.0628$. Both metrics, particularly the correlation dimension, show that the statistical climate of the actual system is effectively reproduced following each transition period, despite the parameter-aware NGRC being trained exclusively on periodic trajectories and is extrapolating the emerging chaotic system behavior. With that the method enables smooth control into unseen target states preventing adverse transient behavior during transitions.
Figure 1: Control method facilitates smooth transitions to unseen target states. The Lorenz system is controlled from a periodic initial state at $\rho_{\text{start}}=100.5$ to a weakly chaotic state at $\rho_{\text{target}}=99.5$, using linearly decreasing parameterizations over various transition durations, shown in D. The control method enables a smooth, transient-free transition to the target state (D, blue). In contrast, simulations with the same parameter transitions (D, red) show that shorter transition durations often cause transients. For each transition time step, 100 differently initialized system were tested. A illustrates one of those initial periodic trajectory at $\rho_{\text{start}}$ (A, red). B and C show the controlled dynamics to the weakly chaotic target state with an instantaneous parameter change (B, blue) and its running correlation dimension (C, green).
@@
Figure 2: Statistical climate of controlled target states. Evaluation of the statistical climate of the $100$ controlled Lorenz systems at $\rho_{target}=99.5$ given each transition period shown in Fig. 1. In A, the correlation dimension is measured and in the B the largest Lyapunov exponent.
Power system model
In this section, we apply the control mechanism on a generic power system model introduced by Dobson et al. [12] that models the system collapse in small power systems (see Methods). We take the reactive power demand $Q_{1}$ as the bifurcation parameter and train the parameter-aware NGRC with 7 differently parametrized datasets in $Q_{1}$ (Fig.3 A). With that we can interpolate and extrapolate the system behavior in unseen parameter regimes and find a periodic window in the vicinity of a system collapse (Fig. 3 D). We take the associated parameter $Q^{window}_{1}=2.989788$ as the target regime $Q^{target}_{1}$ to where we want to control the system to from a periodic state at $Q^{inital}_{1}=2.98940$, far from the control target with rich chaotic behaviors between the parameter range. By analyzing this parameter change through an instantaneous parameter change, using non-stationary simulations with the Runge-Kutta 4 method and 250 different initial conditions, we observe that $35\%$ of the trajectories result in a system collapse (Fig. 3 B), even though the parameterization has not yet reached the critical values typically associated with a system collapse event. Since the system exhibits rich chaotic behaviors between the initial and target parameterizations, $37\%$ of the simulations show trajectories that transition into the periodic target regime with prolonged transient behavior (slow transition). These transitions are characterized by transient periods lasting longer than $12.5$ seconds. The remaining $28\%$ of simulations exhibit fast transitions with transient periods shorter or equal $12.5$ seconds (Fig. 3 C). To quantify the system behavior, we calculated the spectral entropy (see Methods) of the trajectories, which remains constant for periodic dynamics. Since the target dynamic is also periodic, we measure the transient periods as the time required for the spectral entropy to reach a constant value again, adjusted by the overlap used in the short-time fourier transformation domain. The spectral entropy for different system behaviors is shown in Fig. 3 B and E in green and the associated detected transient regions in orange. We apply the control method to control the power system model from the initial periodic regime at $Q^{inital}_{1}$ to the extrapolated periodic window in the vicinity of the system collapse at $Q^{target}_{1}$, under the objective to avoid system collapse and prolonged transient transitions observed for the non-stationary simulations. Therefore, we predict the non-stationary dynamics of the parameter switch from varying initial conditions with the trained NGRC until we find a trajectory that fulfills the control objective. This trajectory is than used to define the control forces applied to each of the 250 previous initial systems, remaining at $Q^{inital}_{1}$ to control them into the target regime at $Q^{target}_{1}$. In Fig. 3 E one of the controlled voltage dynamics with the spectral entropy measure and the detected transient region is displayed. In Fig. 3 F we measured for each of the 250 previous initial systems the transient durations when controlled from their initial parameterization to the target parameterization and find that the control method ensures quantitatively the objective fast transitions, even for initial power system dynamics which would exhibit a system collapse or prolonged transit dynamics. To qualitatively evaluate the controlled dynamics at the target state, we compare their frequency spectrum with that of the actual voltage dynamics at $Q^{target}_{1}$ and define the spectral error as the mean absolute error between the two spectra. In Fig. 4 $\mathbf{A}$ the spectral error is plotted for different control parameters $K$ ranging from 0.25 to 40. For small control parameters this error decreases until a minimum at $K_{min}=4.25$ is reached and then slightly increases until the control method breaks down for control parameters equal or larger than 40, where the dynamics exhibiting a system collapse. The frequency spectrum of the controlled dynamics for $K_{min}$ and the ground truth is plotted in Fig. 4 $\mathbf{B}$, confirming that the NGRC controller is not only able to extrapolate the dynamics and the associated unseen frequencies accurately but also enables the accurate control to these states.
Figure 3: Control method applied to power system control. Fig. A illustrates the voltage behavior of the power system model for several reactive power demands $Q_{1}$. The NGRC is trained on seven training data samples (shown in light green), to accurately interpolate and extrapolate the system behavior in unseen parameter regions of $Q_{1}$ (shown in Fig. D). We test the methods control capabilities by applying an instantaneous parameter switch from an unseen initial regime to an unseen target regime, both of which are periodic (Fig. A cyan). From simulating this switch, three different behaviors occur (Fig. B) with their probability for 250 different initial condition shown in Fig. C. The control method applied to the initial system dynamic, ensures fast transitions to the target dynamic while completely avoid prolonged chaotic transients and system collapses (shown in Fig. E and Fig. F).
Figure 4: Accuracy of the controlled and extrapolated target dynamics. To quantify the accuracy the controlled dynamics in the extrapolated target state, we calculate the mean absolute error of the frequency spectra (Spectral Error) of the simulated and controlled dynamics at the target parametrization. A shows the spectral error of the 250 controlled trajectories as box plots for different control parameters. B shows the mean frequency spectrum of 250 controlled trajectories for the control parameter associated with the smallest spectral error and the target spectrum from simulation.
Power system model with noise
We test the performance of the control method for different levels of additional noise. In this setup the time series of the power system has additional gaussian noise with standard deviation $\sigma$ added to the system at each time step the system is evolving. We test the control objective of fast transitioning from $Q^{initial}_{1}$ to $Q^{target}_{1}$ for different $\sigma$ with the identical controller of the previous results. Only this time the controller processes noisy data $\tilde{\mathbf{u}}(t)=\mathbf{u}(t)+\mathbf{\mu}(t)$ with $\mathbf{\mu}(t)\sim\mathbf{\mathcal{N}}(0,\,\sigma^{2})$ to calculate the control force, which is applied directly onto the (noise free) power system. We measure the transient duration of the actual controlled power system with the spectral entropy. In Fig. 5, we test this setup for different control parameters $K$ and $\sigma$ ranging from 0.002 to 0.05, with 250 initial systems per configuration. While in the noise free case a well chosen control parameter might lead to a slightly reduced spectral error, it becomes increasingly important in the noise case in both the quantitative as well as qualitative metrics. Figure 5 $\mathbf{A}$ shows the fast transition probabilities. The control objective of fast transitioning to the target dynamic is ensured up to $\sigma=0.014$ in a small range of control parameters in the neighborhood of $K=10$ (see Fig. 6 $\mathbf{C}$). For higher noise levels, the fast transition probability drops across all control parameters while the probability of a system collapse generally rises or becomes certain for small and high control parameters in the analyzed range for increasing noise levels. For the noise and control parameter configurations which follow the control objective of ensuring fast transitions, we calculate the spectral error of the frequency spectrum of the controlled target states and the frequency spectrum of the noise free simulated target dynamics in Fig. 5 $\mathbf{C}$. The results show that higher additive noise, which influences the power system though the controller, leads to higher spectral errors. This effect increases with higher control parameters. A comparison of the controlled voltage dynamics in Figs. 6 and 7, featuring different noise levels ($\sigma=0.012$ and $\sigma=0.014$) and control parameter values ($K=30$ and $K=10$), provides a clear visualization of the noise amplification properties of a higher control parameter. The frequency spectra of the dynamics provide the insight that large parts of the observed additional noise is located in the higher frequencies domains above $2Hz$. These domains represent the frequency ranges associated with the system’s time step size, $\Delta t=0.05$, with $f=1/\Delta t=20$, as the control force is applied at each time step, introducing noise to the system. The spectral errors of the control parameters for different noise levels below $2Hz$ are displayed in Fig. 5 $\mathbf{D}$. The results show that in the lower frequency domain the amplification effect of the control parameter nearly vanishes and that the power system is accurately controlled to the frequencies domains of the true target dynamics when applied with noisy input data. To quantify the controller induced noise to the power system we calculate the signal to control noise ratio ($ScNR$) of the controlled voltage dynamics at the target state. We set the voltage of the control dynamic from NGRC as the signal and the differences between the signal and the voltage dynamic of the controlled power system as the control noise. For $K=10$ and $\sigma=0.014$ the $ScNR\approx 14.02$, which means the power of the signal $P_{signal}=10^{14.02/10}P_{nosie}$ is approximately $25$ times larger than the control noise power. For $K=30$ and $\sigma=0.012$, the $ScNR\approx 5.32$, which shows an approximately $3.4$ times larger signal power. In Fig. 8 $\mathbf{B}$ we compare the additive noise levels $\sigma$ with the standard deviation of the control noise $\sigma_{c}$. We find that for small control parameters up to $K=20$, the induced control noises to the power system $\sigma_{c}$ is smaller than the additive noise levels. For $\sigma=0.014$ with $K=10$ the induced $\sigma_{c}$ is approximately $50\%$ smaller. We observe a linear relationships between $\sigma$ in the measured dynamics $\tilde{\mathbf{u}}(t)$ and the control noise $\sigma_{c}$ under different control parameters $K$. The control force $\tilde{\mathbf{F}}(t)=K(\tilde{\mathbf{u}}(t)-\mathbf{v}(t))$ decomposes into $\tilde{\mathbf{F}}(t)=\mathbf{F}(t)+\mathbf{F}_{n}$, where $\mathbf{F}_{n}=K\mathbf{\mu}(t)$. When synchronisation occurs, $\mathbf{F}(t)$ vanishes, leaving only $\mathbf{F}_{n}$. Looking at the standard deviation of $\mathbf{F}_{n}$ and integrating over the time step size $\Delta t$ following Eq. 5, yields $\sigma_{c}=K\Delta t\sigma$, which matches the observed slopes for $\Delta t=0.05$. These results show that for $K>1/\Delta t$ the control noise is amplified, while for smaller control parameters the control noise is damped.
Figure 5: Control with additive noise. The upper plots show the system behavior of the controlled dynamics given different control parameters and noise levels. For each configuration we show the probability with which the 250 controlled dynamics exhibiting the objective fast transition to the target dynamic (A) and a system collapse (B). C shows the spectral error between the fast transitioned controlled target dynamics and the target dynamic without noise. D shows the corresponding error for frequencies below $2Hz$, where the frequencies of the target dynamics are. The results highlight that on the one hand the control method can accurately control the initial dynamics to the target dynamics, even in high noise regimes with a sufficiently high and not higher control parameter $K$. On the other hand choosing an high $K$ leads to additional noise (mainly high frequency) by the control method itself.
Figure 6: Control with additive noise. The upper plots show the voltage dynamics of the power system model with additive gaussian noise $\sigma=0.014$ and its frequency spectrum. The NGRC controller processes the noisy input data and produces a force with control parameter $K=10$ and the differences between its predictions and the noisy input data to control the power system to the target state while maintaining the fast transition period objective (lower plots).
Figure 7: Control with additive noise $\sigma=0.012$. The upper plots show the voltage dynamics of the power system model with additive gaussian noise $\sigma=0.012$ and its frequency spectrum. While the noise is slightly smaller than in Fig. 6, the increased control parameter amplifies the noise in the controlled power system.
Figure 8: Noise induced through control. A displays the mean signal to noise ratios for the $250$ fast transitioning controlled dynamics at the target dynamic with their standard deviation for different control parameters $K$ and additive noise levels $\sigma$. B shows the control noise on the controlled target dynamics originating from the control method itself for different additive noise levels and control parameters.